For the fusion reaction shown, calculate the change in energy of the reaction in units of joules per mole. H12+H12⟶H13+H11

Atom or particle Mass (amu)

H-1 1.00782

H-2 2.01410

H-3 3.01605

He-3 3.01603

He-4 4.00260

neutron 1.00866
Answer #1

lets calculate the mass of all the reactant

mass of reactant

= 2*mass(H12)

=2*2.0141

= 4.0282

lets calculate the mass of all the product

mass of product

= 1*mass(H13) + 1*mass(H11)

=1*3.01605 + 1*1.00782

= 4.02387

Decrease in mass,

dm = net mass of reactant - net mass of product

= 4.0282 - 4.02387

= 0.00433

This decrease in mass will be converted to energy

dm = 4.33*10^-3 u

1 u = 1.6605*10^-27 Kg

So,

dm = 4.33*10^-3*1.6605*10^-27

= 7.19*10^-30 Kg

mass can be converted to energy as given by:

E = m*c^2

so,

E = 7.19*10^-30*(3*10^8)^2 J

E = 6.471*10^-13 J

This is energy of 1 atom

Energy of 1 mol = energy of 1 atom * Avogadro's number

= 6.471*10^-13 *6.022*10^23 J/mol

= 3.897*10^11 J/mol

Answer: 3.9*10^11 J/mol